# Broken Expectations: The Effects of Modelling Assumptions on the Inferred Dark Matter Distribution in the Milky Way’s Satellites

### v1

@@ Appendix B Velocity anisotropy recovery
Appendix C Effect of data truncation
In this Section, we investigate in more detail the quality of the pyGravSphere models of Fornax. Additionally, we test whether truncation of the sample stellar particles fed into pyGravSphere makes a difference in the quality of the recovered density profiles, motivated by the fact that the line-of-sight velocity dispersion profiles that we analyse in this work extend much further than the currently available kinematic data (from which a velocity dispersion profile may be resolved).
The top panel in Fig. C1, shows the recovery by pyGravSphere of the density profile of Fornax in the ‘light’ potential, with the case of full data (top row) and data truncated at 2 kpc (bottom row). The anisotropy profile is not recovered well in either case. When the data is truncated, pyGravSphere prefers more tangential anisotropies in the outskirts, which worsens the quality of the recovered density profile in the outer parts of the galaxy. From the fit to the line-of-sight velocity dispersion, it is clear that the truncation has resulted in pyGravSphere preferring models with declining velocity dispersion beyond 2 kpc since the available data does not fully capture the rise in the velocity dispersion profile. It is also clear that the virial shape parameters disagree between the two data sets – most importantly, the “true” VSPs, which are estimated from the data. This indicates that pyGravSphere’s method of evaluating the “true” virial shape parameters can be biased, especially when the data is close to complete. In practice, the estimation of the “true” VSPs involves fitting a power law to the fourth moment of velocity profile outside of the half-light radius and further marginalising over how far the profile extends (which is taken to be up to 2 times the outermost data point) and how steeply it falls off beyond that point (Read & Steger, 2017; Read et al., 2018). For our simulated Fornax in the ‘light’ potential, where the fourth velocity moment is rising in the outskirts of the galaxy, for the case with full data, this procedure is likely to overestimate the virial shape parameters because the true value is on the lower edge of the area marginalised over. For spatially truncated data, this is likely to underestimate the VSPs, because full data extends further than twice the half-light radius. This is demonstrated in the top panel of Fig. C2.
For the ‘heavy’ Fornax, where the velocity dispersion profile is declining, truncating the data at 2 kpc has still captured the decline of the profile beyond the ‘bump’ at 1.5 kpc. The estimated ‘true’ VSPs are in agreement with each other in both cases. This is because the 4th moment of the velocity profile declines in the full-data case in the outer parts of the galaxy, while in the truncated case, it is fairly flat. Thus, the estimate of the virial shape parameters is slightly higher in the truncated-data case. This is demonstrated in the bottom panel of Fig C2. The noteworthy difference is, instead, the shape of the recovered velocity dispersion profile in the inner regions, which declines towards the smaller radii in the full-data case, but is flat when the data is truncated. This is reflected in the recovery of the correct isotropic orbits in the centre. The best-fitting Jeans models have VSP2 km4s-4, permitted by the larger error bars on the “true” VSP2 estimate. It is thus not fully clear why this model performs better, given that the best-fitting Jeans model VSP2 in the truncated-data case is in disagreement with the true one obtained with full data, while VSP1 is nearly identical between the two models. The bottom panel of Fig. C2 demonstrates that this high value of VSP2 can only be achieved by a shallower tail of the fourth-moment profile than is allowed by the full data. This suggests that either the better fit to the inner density profile in the truncated data case is some peculiarity of the employed models, or the declining tail in the fourth moment, that is a result of enhanced tidal forces in the ‘heavy’ Milky Way potential, is unusable in the dynamical modelling. We discuss this more in the next section.
Given our results, we suggest that in the future applications of pyGravSphere, the virial shape parameters are not estimated for radii where there is no data as in Read et al. (2018), but instead computed only for the extent of the data and used as lower limits on pyGravSphere models.
Appendix D Power Law Smoothness Constraint
The parameter , which controls the maximum difference between the slopes of two neighbouring bins, is set to 1 in pyGravSphere. This ensures that the returned models have smoothly decreasing densities, as predicted by cosmological -body simulations. We explore the effect of this constraint here, with and without the virial shape parameters.
In the first column of Fig. D3, we show the results of the power-law model without virial shape parameters applied to our simulated Fornax in the ‘light’ and ‘heavy’ Milky Way potential. The ‘true’ density profiles are recovered within , while in the ‘heavy’ case, the models underestimate the density in the outer radii. When the virial shape parameters are employed (second column), there is a visibly stronger constraint on both the inner and outer profiles; however, the underestimation of the central density and overestimation near the half-light radius that we have seen throughout this work becomes more prominent. Evidently, the inclusion of virial shape parameters has favoured models with shallower inner density slopes.
In the third column, we now remove the parameter from the modelling, allowing each power-law bin to vary freely within its priors. The recovered density profiles now strongly prefer a more cuspy solution that is in excellent agreement with the truth. When the virial shape parameters are employed (fourth column) the recovered profile remains in excellent agreement with the truth in the inner regions, while the power-law bin between 1-2 is clearly biased towards shallower slopes in both the ‘light’ and ‘heavy’ cases. Upon inspecting all of our models, including ones with the ‘smoothness’ constraint, it seems that there is a preference for such a feature in the profile for Fornax. When the ‘smoothness’ constraint is applied, the power-law slopes below that would necessarily need to be shallower, thus favouring more core-like inner densities. This effect is exaggerated when the VSPs are employed, as the range of slopes for this power-law bin is significantly reduced.
It is not entirely clear what the origin of this preferred shallower slope near the half-light radius is, but we suspect it may be associated with the turnover features of the line-of-sight velocity dispersion profiles and how the data is binned in these regions. We plan to investigate this in future work.

### v2

@@ Appendix B Velocity anisotropy recovery
Appendix C Choice of Power law break radii
In the original formulation of the GravSphere model (Read & Steger, 2017), the breaks in the broken-power-law density profile are defined as multiples of the projected half-light radius , motivated by the mass content being most strongly constrained within . Potential dark matter core formation due to heating is also expected to occur within approximately the half-light radius of the dwarf, motivating additional binning in of the profile in this region. Read & Steger (2017) considered mock data with as much as stellar tracers and found that the choice of dark matter binning did not alter their results significantly. However, this result ultimately depends on how embedded the stellar profile is inside the dark matter halo and whether the multiples of can capture the major changes in the log-slope of the dark matter profile, particularly near the dark matter scale radius and the tidal radius. Compared to the mocks discussed in detail in Read & Steger (2017) and Read et al. (2021), the half-light radii of our dwarfs are significantly more embedded within the dark matter scale radius (typically within 10 per cent of the dark matter scale radius), meaning that 2-4 does not capture the region beyond the dark matter scale radius. When the dwarfs are tidally stripped, the dark matter density profile is additionally steepened. The accuracy of the broken power-law fit therefore depends on whether the region 2-4 can capture the significant change in the log-slope of the dark matter. In fact, a good predictor of GravSphere’s performance on mocks analysed in Read et al. (2021) is how embedded the stellar half-light radius is within the dark matter profile scale radius, with the most inaccurate results arising from the cases where . The availability of kinematic tracers at these radii likely has a secondary effect, whereby the virial shape parameters put a stronger constraint on the outer slope, which affects the recovery of the inner profile via the smoothness criterion.
In Fig. C3 we investigate this by directly fitting the power law model to the known density profile of the Fornax ‘light’ model at the present day, without the inclusion of the smoothness criterion. The ‘kick’ in the profile near 1 kpc can be seen, which is more subtle in Fig. 8 (see also the Supplementary Materials). We then test this result by splitting the profile into regions that are fit by individual power laws. The first four power laws give an excellent fit to the profile, but the fifth power law () attempts to fit the outermost steeper regions of the halo. It is straightforward to see how this results in the shape of the overall fit (black line) and how this would lead to inferred core-like profiles in the inner regions if slopes below have to be shallower. In Supplementary Materials, we also show that applying the GravSphere model without the smoothness criterion leads to a recovered profile similar to the black line in Fig. C3.
This problem can of course be solved by adding additional bins to the power-law model. Ideally, a range of power-law bins can be tested to ensure convergence of the result. Bañares-Hernández et al. (2025) have also utilised our Fornax ‘light’ model to test the GravSphere2 code that utilises the coreNFWtides dark matter model and achieved a more accurate dark matter profile recovery. Here, we also test a simple modification to the broken-power-law model, where the break radius is treated as a free parameter, instead of being set to the projected half-light radius. As shown in Fig. C4, this leads to an accurate recovery of the profile in the outer parts, while, at the centre, a more cuspy solution is preferred, since 0.25 times the best-fitting corresponds to approximately 0.5 kpc and cannot fully capture the inner regions of the profile.
This explains in part why the effect is less significant in Draco and Ursa Minor, since the data does not reach far enough to capture kinematic features associated with the change in dark matter density. As we show in the Supplementary Materials, data truncation to the central 2 kpc has resulted in more accurate recovery of the dark matter density in the Fornax ‘light’ and ‘heavy’ model. However, one has to bear in mind that the projected central regions of the dwarf also carry information about the outer density profile, and thus the bias cannot be avoided fully with data truncation when the model is a poor representation of the true dark matter density profile.